# Efficient and Generalizable Certified Unlearning: A Hessian-free Recollection Approach

### v3

@@ B.4 Detailed Proof of Theorem 4
$\|\mathbf{w}_{E,B}^{-u}$ $-\mathbf{w}_{E,B}-\mathbf{a}_{E,B}^{-u}\|$ (41)
$\leq$ $\ \ \eta\bigg{(}q^{T-1}\|\Delta^{-u}_{E,B-1}\|+q^{T-2}\|\hat{\mathbf{H}}_{E,B-1}\Delta^{-u}_{E,B-2}\|+...+q^{0}\prod^{E}_{k=0}\prod_{b=0}^{B-1}\hat{\mathbf{H}}_{k,b}\Delta^{-u}_{0,0}\bigg{)}$
@@
$\leq$ $\ \ 2\eta^{2}C\frac{q^{T-1}-q^{2(T-1)}}{1-q}+2\eta^{2}C\frac{q^{T-2}-q^{2(T-2)}}{1-q}\rho+...+2\eta^{2}C\frac{q^{0}-q^{2\times(0)}}{1-q}\rho^{T-1}$
$\ \ +\frac{2C\eta}{|\mathcal{B}|}q^{T-B+b(u)}\rho^{B-b(u)-1}+...+\frac{2C\eta}{|\mathcal{B}|}q^{b(u)}\rho^{T-b(u)-1}$
$=$ $\ \ \frac{2\eta^{2}C}{1-q}\sum_{k=0}^{T-1}\rho^{T-k-1}(q^{k}-q^{2k})+\frac{2C\eta}{|\mathcal{B}|}\sum_{e=0}^{E}\rho^{T-eB-b(u)-1}q^{eB+b(u)}.$
Through the polynomial multiplies geometric progression, we obtain,
$\|\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}-\mathbf{a}_{E,B}^{-u}\|$ $=\frac{2\eta^{2}C}{1-q}\sum_{k=0}^{T-1}\rho^{T-k-1}(q^{k}-q^{2k})+\frac{2C\eta}{|\mathcal{B}|}\sum_{e=0}^{E}\rho^{T-eB-b(u)-1}q^{eB+b(u)}$ (42)
$=\frac{2\eta^{2}C}{1-q}\rho^{T-1}\sum_{k=0}^{T-1}\left((\frac{q}{\rho})^{k}-(\frac{q}{\rho}^{2})^{k}\right)+\frac{2C\eta}{|\mathcal{B}|}\rho^{T-b(u)-1}q^{b(u)}\sum_{e=0}^{E}(\frac{q}{\rho})^{eB}$
$=\frac{2\eta^{2}C}{1-q}\left(\frac{\rho^{T}-q^{T}}{\rho-q}-\frac{\rho^{T}-q^{2T}}{\rho-q^{2}}\right)+\frac{2C\eta}{|\mathcal{B}|}\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}}\rho^{B-b(u)-1}q^{b(u)}.$
For convenience, we define $\zeta_{T}^{-u}=\frac{\eta}{1-q}(\frac{\rho^{T}-q^{T}}{\rho-q}-\frac{\rho^{T}-q^{2T}}{\rho-q^{2}})+\frac{1}{|\mathcal{B}|}\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}}\rho^{B-b(u)-1}q^{b(u)}$. At this point, we have completed the proof of Theorem 4 as follows,
$\|\mathbf{w}_{E,B}^{-u}-\mathbf{w}_{E,B}-\mathbf{a}_{E,B}^{-u}\|\leq 2\eta C\zeta_{T}^{-u}.$ (43)
Note that $2\eta C$ is a small constant. Therefore, when $\rho<1$ (e.g. Assumption 1), where $T$ is sufficiently large, the error tends towards zero. That is, the unlearned model can almost be equivalent to the retrained model without the need for retraining. ∎
B.5 Additional Definitions
@@
The loss function $\ell(\mathbf{w};z)$ is $\lambda$-strongly convex, i.e., there exists a constant $\lambda>0$ such that for any $z\in\mathcal{Z}$ and $\mathbf{w}_{1},\ \mathbf{w}_{2}\in\mathbb{R}^{d}$, it holds that
@@
or equivalently, $\left\|\nabla\ell^{2}\left(\mathbf{w};z\right)\right\|\geq\lambda$ for all $\mathbf{w}\in\mathbb{R}^{d}$.
@@
The loss function $\ell(\mathbf{w};z)$ is $L$-Lipschitz, i.e., there exists a constant $L>0$ such that for any $z\in\mathcal{Z}$ and $\mathbf{w}_{1},\ \mathbf{w}_{2}\in\mathbb{R}^{d}$, it holds that
@@
Definition 4 ($M$-Smoothness or gradient-Lipschitzness).
The loss function $\ell(\mathbf{w};z)$ is $M$-smooth, i.e., there exists a constant $M>0$ such that for any $z\in\mathcal{Z}$ and $\mathbf{w}_{1},\ \mathbf{w}_{2}\in\mathbb{R}^{d}$, it holds that
@@
or equivalently, $\left\|\nabla\ell^{2}\left(\mathbf{w};z\right)\right\|\leq M$ for all $\mathbf{w}\in\mathbb{R}^{d}$.
B.6 Detailed Proof of Theorem 5
@@
Lemma 6 (Shalev-Shwartz et al. [46]).
Let $\mathbf{w}^{*}$ denote a minimizer of the population risk in (1) and $\hat{\mathbf{w}}^{-U}$ denote a minimizer of the empirical risk without the knowledge of $U$ in (2) which aims to minimize $F_{\mathcal{S}\backslash U}(\mathbf{w})$, where $|U|=m$. For $\lambda$-strongly convex and $L$-lipschitz objectives, we have,
@@
A comprehensive proof of Lemma 6 is available in [46]. ∎
@@
For any $z\in\mathcal{Z}$, the loss function $\ell(\mathbf{w};z)$ is $\lambda$-strongly convex, $L$-Lipschitz and $M$-smooth. Let $\hat{\mathbf{w}}^{-U}=\operatorname{argmin}_{\mathbf{w}}F_{\mathcal{S}\backslash U}(\mathbf{w})$ without the knowledge of $U$ in problem (2). For all $e,b$ and vector $\mathbf{v}\in\mathbb{R}^{d}$, the spectral radius of $\mathbf{I}-\frac{\eta_{e,b}}{|\mathcal{B}_{e,b}|}\mathbf{H}_{e,b}$ is defined as $\rho$ which is largest absolute eigenvalue of these matrices. We have that after $T$ steps of GD with initial stepsize $\eta\leq\frac{2}{\lambda+M}$,
@@
Let $\eta\leq\frac{2}{\lambda+M}$. Therefore, $\eta_{T}\leq\frac{2}{\lambda+M}$, which implies that for any $T\geq 0$:
@@
Let $\hat{\mathbf{w}}^{-U}$ be the empirical risk minimizer of the objective function $F_{\mathcal{S}\backslash U}(\mathbf{w})$, where $|U|=m$. Let $T$ be the total number of update steps. For any $z\in\mathcal{Z}$ the loss function $\ell(\mathbf{w};z)$ is $\lambda$-strongly convex, $M$-smoothness and $L$-Lipschitz. For a minimizer $\mathbf{w}^{*}$ of the population risk in (1) and $\tilde{\mathbf{w}}_{E,B}^{-U}=\mathbf{w}_{E,B}+\mathbf{a}_{E,B}^{-U}+\sigma$, we have that
@@
For the first term of (58), we have
@@
Therefore, we now bound the first term $C_{3}$ and the second term $C_{4}$ of (59) as follows:
@@
$\mathbb{E}\big{[}\|{\mathbf{w}}_{E,B}^{-U}-\hat{\mathbf{w}}^{-U}\|\big{]}\leq\rho^{T}\frac{2L}{\lambda},$ (60)
where $\text{where }\rho$ $=\max\left\{\left|1-\eta\lambda\right|,\left|1-\eta M\right|\right\}<1.$
@@
$\mathbb{E}[\|\tilde{\mathbf{w}}_{E,B}^{-U}-{\mathbf{w}}_{E,B}^{-U}\|]$ $=\mathbb{E}[\|\mathbf{w}_{E,B}+\mathbf{a}_{E,B}^{-U}-\mathbf{w}_{E,B}^{-U}\|]+\mathbb{E}[\|\sigma\|]$ (61)
$\leq 2\eta C\zeta_{T}^{-U}+\sqrt{d}c.$
$\text{where}\ \zeta_{T}^{-U}$ $=\mathcal{O}(\frac{\rho^{T}-q^{T}}{\rho-q}-\frac{\rho^{T}-q^{2T}}{\rho-q^{2}}+\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}}).$
And since the loss function $\ell(\mathbf{w};z)$ satisfies $M$-smoothness and $\lambda$-strongly convex, we thus have that $10<\rho<$ 1. Plugging the above bound (60) and (61) into (59), we have
$\mathbb{E}\left[F(\tilde{\mathbf{w}}_{E,B}^{-U})-F\left(\mathbf{w}^{*}\right)\right]\leq\rho^{T}\frac{2L^{2}}{\lambda}+(\frac{\sqrt{d}\sqrt{2\ln{(1.25/\delta)}}}{\epsilon}+1)\cdot{2L\eta C}\zeta_{T}^{-U}+\frac{4L^{2}}{\lambda(n-m)}.$ (62)
@@
Tightness Analysis of Theorem 5. (1) When $T$ is sufficiently large such that the model $\mathbf{w}_{E,B}^{-U}$ is the empirical risk minimizer $\hat{\mathbf{w}}^{-U}$ of the objective function $F_{\mathcal{S}\backslash U}(\mathbf{w})$, the first two terms tend towards zero: the first term $\|{\mathbf{w}}_{E,B}^{-U}-\hat{\mathbf{w}}^{-U}\|$ is zero; In the second term, we observe that $\zeta_{T}^{-U}=\mathcal{O}(\frac{\rho^{T}-q^{T}}{\rho-q}-\frac{\rho^{T}-q^{2T}}{\rho-q^{2}}+\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}})$, and as $T$ is sufficiently large, the error tends towards zero.
@@
(2) Furthermore,the following settings, where the model does not converge and $\rho$ is close to $q$, might render the bound on the approximation error vacuous and also lead to high approximation error. We note that the first and third terms of $\zeta_{T}^{-U}$ can be equivalently written as $\frac{\rho^{T}-q^{T}}{\rho-q}+\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}}=\rho^{T-1}\sum_{t=0}^{T-1}(\frac{q}{\rho})^{t}+\rho^{T-B}\sum_{e=0}^{B}(\frac{q}{\rho})^{eB}$. These terms increase as $\rho$ approaches $q$, however are always upper bounded by $\rho^{T-1}\cdot T+\rho^{T-B}\cdot T$. In our analysis, $T=EB+B$ is a fixed constant, so this upper bound remains meaningful, although it may not be tight in some cases.
Appendix C Supplementary Experiments
@@
The experiments were conducted on the NVIDIA GeForce RTX 4090 (3090). The code were implemented in PyTorch 2.0.0 and leverage the CUDA Toolkit version 11.8. Our comprehensive tests were conducted on AMD EPYC 7763 CPU @1.50GHz with 64 cores under Ubuntu20.04.6 LTS.
Source Code: https://github.com/Anonymous202401/If-Recollecting-were-Forgetting
C.1 Additional Verification Experiments
@@
The verification experiments in this section consist of two parts: (1) We first investigate the reasons behind the randomness of previous works under the non-convex setting. We select forgetting samples under different random seeds, and show that previous works exhibit dependence for forgetting samples, meaning that the forgetting effect of previous works is beneficial for some samples but poor for others. (2) Secondly, we conduct experiments on larger-scale dataset and model with more metrics, which previous works could not achieve because the computation of the Hessian is expensive. This demonstrates the potential of our proposed hessian-free method on large-scale models and datasets.
@@
Sensitivity to the forgetting samples. We conduct experiments on a simple CNN and the hyperparameter setup is consistent with the verification experiments I in the main paper. As shown in Figure 2, we provide intuitive insights into the reasons for the decline of prior works in non-convex settings. Specifically, we investigated the $L_{2}$ norm of the approximate parameter change $\|\mathbf{a}^{-U}\|$ and the norm of the exact parameter change $\|\mathbf{w}_{E,B}^{-U}-\mathbf{w}_{E,B}\|$ across different selections of the forgetting sample. We observed that the performance of NS and IJ depends on the selection of the forgetting data, i.e., they can approximate actual changes well for some data points while generating large approximation errors for others under the non-convex setting, as illustrated in Figure 2. The dependency of forgetting data results in the decrease in distance and correlation coefficient metrics observed in Figure 1 (b). The observed results for NS and IJ are consistent with previous studies on influence functions [35, 42], where these methods display random behavior under the non-convex setting, resulting in a decrease in correlation with retraining. In contrast, our approach can effectively approximate the actual norm of parameter changes, regardless of the forgetting sample selection.
@@
Figure 2: Verification experiments II. Evaluation shows a comparison comparison between the norm of approximate parameter change $\|\mathbf{a}^{-U}\|$ and norm of exact parameter change $\|\mathbf{w}_{E,B}^{-U}-\mathbf{w}_{E,B}\|$ across different random seeds. Intuitively, the NS and IJ methods are contingent on the selection of forgetting data. In contrast, our approach consistently approximates the actual values effectively.
@@
Additional Experiments. We further evaluate using a larger-scale model (ResNet-18 [47]) which features 11 M parameters with three different datasets: CIFAR-10 [48] for image classification, CelebA [49] for gender prediction, and the LFWPeople [50] for face recognition across 29 different individuals. Since prior second-order works NS [2] and IJ [3] are difficult to compute (Hessian) in this scenario, we instead use the following baselines, as described and similarly set up in [25]:
@@
  • FineTune: In case of finetuning, the original learned model is finetuned on the remaining dataset.
@@
  • NegGrad: In case of gradient ascent, the learned model is finetuned using the negative of update direction on the forgetting dataset.
@@
Configurations: (1) We conducted evaluation on ResNet-18 trained on CIFAR-10 with 50,000 samples. The learning stage is conducted for 40 epochs with a stepsize of 0.001 and a batch size of 256. For FineTune, 10 epochs of training are performed with a step size of 0.001 and a batch size of 256. For NegGrad, we run gradient ascent for 1 epoch with a step size of $5\times 10^{-5}$ and a batch size of 2. (2) We conducted evaluation on ResNet-18 trained on LFW with 984 samples, for the classification of 29 facial categories. The learning stage is conducted for 49 epochs with a stepsize of 0.004 and a batch size of 41. For FineTune, 10 epochs of training are performed with a step size of 0.004 and a batch size of 41. For NegGrad, we run gradient ascent for 1 epochs with a step size of $4\times 10^{-6}$ and a batch size of 1. (3) We conducted evaluation on ResNet-18 trained on CelebA with 10,000 samples. The learning stage is conducted for 5 epochs with a stepsize of 0.01 and a batch size of 64. For FineTune, 2 epochs of training are performed with a step size of 0.01 and a batch size of 64. For NegGrad, we run gradient ascent for 1 epoch with a step size of 0.0005 and a batch size of 1. Finally, we uniformly unlearned 50 data points for each dataset. We define $D_{t}$ as the test dataset, $D_{r}$ as the remaining dataset, and $D_{f}$ as the forgetting dataset.
@@
The following Table 3-5 are our evaluation results: we employ five widely-used metrics to evaluate the performance of unlearning methods: Acc_$D_{t}$ (accuracy on $D_{t}$), error _$D_{t}$ (error on $D_{t}$), error_$D_{r}$ (error on $D_{r}$), error_$D_{f}$ (error on $D_{f}$), dis. ($L_{2}$ distance between the unlearned and retrained model).
@@
Performances of our Scheme: We still maintain model indistinguishability and accuracy indistinguishability from the retrained model on more complex tasks and larger models, where the distance is only 0.06 on CIFAR10 and 0.29 on CelebA. The results also shows that the unlearned model of proposed method performs similarly to the retrained model in terms of the accuracy metric. Moreover, compared to these baselines, our method achieves forgetting in only millisecond-level due to the pre-computation of all statistical data, demonstrating the potential of our method on larger-scale models. Additionally, we noticed that when training is unstable, our method would lead to larger performance degradation, such as on the LFW dataset. Although our method has similar accuracy on the forgetting dataset, it suffers performance loss on the remaining dataset. This implies that our method is sensitive to the training process and hyperparameters in non-convex scenarios. We summarize the limitations in Subsection E.2, and provide possible schemes to address these problems.

### v4

@@ C.4 Detailed Proof of Theorem 4
$\|\mathbf{w}_{E,B}^{-u}-\bar{\mathbf{w}}_{E,B}^{-u}\|$ $\leq\eta\left(q^{T-1}\|\Delta\mathbf{w}^{-u}_{E,B-1}\|+q^{T-2}\|\hat{\mathbf{H}}_{E,B-1}\Delta\mathbf{w}^{-u}_{E,B-2}\|+...+q^{0}\prod^{E}_{k=0}\prod_{b=0}^{B-1}\hat{\mathbf{H}}_{k,b}\Delta\mathbf{w}^{-u}_{0,0}\right)$ (55)
@@
$\leq$ $\ \ 2\eta^{2}G\frac{q^{T-1}-q^{2(T-1)}}{1-q}+2\eta^{2}G\frac{q^{T-2}-q^{2(T-2)}}{1-q}\rho+...+2\eta^{2}C\frac{q^{0}-q^{2\times(0)}}{1-q}\rho^{T-1}$
$\ \ +\frac{2\eta G}{|\mathcal{B}|}q^{T-B+b(u)}\rho^{B-b(u)-1}+...+\frac{2G\eta}{|\mathcal{B}|}q^{b(u)}\rho^{T-b(u)-1}$
$=$ $\ \ \frac{2\eta^{2}G}{1-q}\sum_{k=0}^{T-1}\rho^{T-k-1}(q^{k}-q^{2k})+\frac{2\eta G}{|\mathcal{B}|}\sum_{e=0}^{E}\rho^{T-eB-b(u)-1}q^{eB+b(u)}.$
Suppose $q<\rho$. Through the polynomial multiplies geometric progression, we obtain,
$\|\mathbf{w}_{E,B}^{-u}-\bar{\mathbf{w}}_{E,B}^{-u}\|$ $=\frac{2\eta^{2}G}{1-q}\rho^{T-1}\sum_{k=0}^{T-1}\left((\frac{q}{\rho})^{k}-(\frac{q}{\rho}^{2})^{k}\right)+\frac{2\eta G}{|\mathcal{B}|}\rho^{T-b(u)-1}q^{b(u)}\sum_{e=0}^{E}(\frac{q}{\rho})^{eB}$ (56)
$=\frac{2\eta^{2}G}{1-q}\left(\frac{\rho^{T}-q^{T}}{\rho-q}-\frac{\rho^{T}-q^{2T}}{\rho-q^{2}}\right)+\frac{2\eta G}{|\mathcal{B}|}\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}}\rho^{B-b(u)-1}q^{b(u)},$
which yields an upper bound on the approximation error. For ease of comparison, we subsequently simplify the aforementioned result, as follows:
$\|\mathbf{w}_{E,B}^{-u}-\bar{\mathbf{w}}_{E,B}^{-u}\|$ $\leq\frac{2\eta^{2}G}{1-q}\left(\frac{\rho^{T}-q^{2T}}{\rho-q}-\frac{\rho^{T}-q^{2T}}{\rho-q^{2}}\right)+\frac{2\eta G}{|\mathcal{B}|}\frac{\rho^{T}-q^{2T}}{\rho^{B}-q^{B}}\rho^{B-1}$ (57)
$\leq 2\eta G(\rho^{T}-q^{2T})\left(\frac{\eta q}{(\rho-q)(\rho-q^{2})}+\frac{1}{|\mathcal{B}|(\rho-q)}\right).$
At this point, we have completed the proof of Theorem 4 as follows,
$\|\mathbf{w}_{E,B}^{-u}-\bar{\mathbf{w}}_{E,B}^{-u}\|\leq\mathcal{O}(2\eta G\rho^{T}).$ (58)
If Theorem 4 further satisfies Assumption 1, and we choose $\eta\leq{2}/{(M+\lambda)}$ to ensure gradient descent convergence as well as $\rho<1$, we thus obtain a tight upper bound of $\mathcal{O}(\rho^{n})$ for each sample. ∎
C.5 Additional Definitions of Assumption 1
@@
The loss function $\ell(\mathbf{w};z)$ is $\lambda$-strongly convex with respect to $\mathbf{w}$, i.e., there exists a constant $\lambda>0$ such that for any $z\in\mathcal{Z}$ and $\mathbf{w}_{1},\ \mathbf{w}_{2}\in\mathbb{R}^{d}$, it holds that
@@
or equivalently, $\left\|\nabla^{2}\ell\left(\mathbf{w};z\right)\right\|\geq\lambda$ for all $\mathbf{w}\in\mathbb{R}^{d}$.
@@
The loss function $\ell(\mathbf{w};z)$ is $L$-Lipschitz with respect to $\mathbf{w}$, i.e., there exists a constant $L>0$ such that for any $z\in\mathcal{Z}$ and $\mathbf{w}_{1},\ \mathbf{w}_{2}\in\mathbb{R}^{d}$, it holds that
@@
Definition 5 ($M$-Smoothness).
The loss function $\ell(\mathbf{w};z)$ is $M$-smooth with respect to $\mathbf{w}$, i.e., there exists a constant $M>0$ such that for any $z\in\mathcal{Z}$ and $\mathbf{w}_{1},\ \mathbf{w}_{2}\in\mathbb{R}^{d}$, it holds that
@@
or equivalently, $\left\|\nabla^{2}\ell\left(\mathbf{w};z\right)\right\|\leq M$ for all $\mathbf{w}\in\mathbb{R}^{d}$.
C.6 Detailed Proof of Theorem 6
@@
Lemma 9 (Shalev-Shwartz et al. (2009)).
Let $\mathcal{S}=\left\{z_{i}\right\}_{i=1}^{n}$ where $S\sim\mathcal{D}^{n}$. Let $\mathbf{w}^{*}$ denote a minimizer of the population risk in Eq. 1 and $\hat{\mathbf{w}}^{-U}$ denote a minimizer of the empirical risk without the knowledge of $U$ in Eq. 2 which aims to minimize $F_{\mathcal{S}\backslash U}(\mathbf{w})$, where $|U|=m$. Let Assumption 1 hold, we have,
@@
A comprehensive proof of Lemma 9 is available in Shalev-Shwartz et al. (2009). ∎
For brevity and to save space, we present our generalization guarantee only for GD. Nevertheless, the results for SGD can be analogously obtained by similar process.
@@
For any $z\in\mathcal{Z}$, the loss function $\ell(\mathbf{w};z)$ is $\lambda$-strongly convex, $L$-Lipschitz and $M$-smooth. Let $\hat{\mathbf{w}}^{-U}=\operatorname{argmin}_{\mathbf{w}}F_{\mathcal{S}\backslash U}(\mathbf{w})$ without the knowledge of $U$ in problem Eq. 2. For all $e,b$ and vector $\mathbf{v}\in\mathbb{R}^{d}$, the spectral radius of $\mathbf{I}-{\eta_{e,b}}\mathbf{H}_{e,b}$ is defined as $\rho$ which is largest absolute eigenvalue of these matrices. We have that after $T$ steps of GD with initial stepsize $\eta\leq\frac{2}{\lambda+M}$,
@@
Let $\eta\leq\frac{2}{\lambda+M}$. Therefore, $\eta_{T}\leq\frac{2}{\lambda+M}$, which implies that for any $T\in[0,\infty)$:
@@
Let $\hat{\mathbf{w}}^{-U}$ be the empirical risk minimizer of the objective function $F_{\mathcal{S}\backslash U}(\mathbf{w})$, where $|U|=m$. Let $T$ be the total number of update steps. For any $z\in\mathcal{Z}$ the loss function $\ell(\mathbf{w};z)$ is $\lambda$-strongly convex, $M$-smoothness and $L$-Lipschitz. For a minimizer $\mathbf{w}^{*}$ of the population risk in Eq. 1 and $\tilde{\mathbf{w}}_{E,B}^{-U}=\mathbf{w}_{E,B}+\mathbf{a}_{E,B}^{-U}+\mathbf{n}$, we have that
@@
For the first term of Eq. 73, we have
@@
Therefore, we now bound the first term $C_{3}$ and the second term $C_{4}$ of Eq. 74 as follows:
@@
$\mathbb{E}\big{[}\|{\mathbf{w}}_{E,B}^{-U}-\hat{\mathbf{w}}^{-U}\|\big{]}\leq\rho^{T}\frac{2L}{\lambda},\text{where }\rho=\max\left\{\left|1-\eta\lambda\right|,\left|1-\eta M\right|\right\}<1.$ (75)
@@
$\mathbb{E}[\|\tilde{\mathbf{w}}_{E,B}^{-U}-{\mathbf{w}}_{E,B}^{-U}\|]$ $=\mathbb{E}[\|\mathbf{w}_{E,B}+\mathbf{a}_{E,B}^{-U}-\mathbf{w}_{E,B}^{-U}\|]+\mathbb{E}[\|\mathbf{n}\|]\leq\|\Delta_{E,B}^{-U}\|+\sqrt{d}c.$ (76)
$\text{where}\ \|\Delta_{E,B}^{-U}\|$ $=\frac{2\eta^{2}G}{1-q}\left(\frac{\rho^{T}-q^{T}}{\rho-q}-\frac{\rho^{T}-q^{2T}}{\rho-q^{2}}\right)+\frac{2\eta G}{|\mathcal{B}|}\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}}\rho^{B-b(u)-1}q^{b(u)}.$
And since the loss function $\ell(\mathbf{w};z)$ satisfies $M$-smoothness and $\lambda$-strongly convex, we thus have that $10<\rho<$ 1. Plugging the above bound Eq. 75 and Eq. 76 into Eq. 74, we have
$\mathbb{E}\left[F(\tilde{\mathbf{w}}_{E,B}^{-U})-F\left(\mathbf{w}^{*}\right)\right]\leq\frac{4L^{2}}{\lambda(n-m)}+\rho^{T}\frac{2L^{2}}{\lambda}+(\frac{\sqrt{d}\sqrt{2\ln{(1.25/\delta)}}}{\epsilon}+1)\cdot{L\|\Delta_{E,B}^{-U}\|}.$ (77)
@@
Tightness Analysis of Theorem 6. Furthermore, the following settings, where $\rho$ is close to $q$, might render the bound on the approximation error vacuous and also lead to a high approximation error. We note that the first and third terms of $\|\Delta_{E,B}^{-U}\|$ can be equivalently written as $\mathcal{O}(\frac{\rho^{T}-q^{T}}{\rho-q}+\frac{\rho^{T}-q^{T}}{\rho^{B}-q^{B}}=\rho^{T-1}\sum_{t=0}^{T-1}(\frac{q}{\rho})^{t}+\rho^{T-B}\sum_{e=0}^{B}(\frac{q}{\rho})^{eB})$. These terms increase as $\rho$ approaches $q$, however are always upper bounded by $\mathcal{O}(\rho^{T-1}\cdot T+\rho^{T-B}\cdot T)$. In our analysis, $T=EB+B$ is a fixed constant, so this bound remains meaningful, although it may not be tight in some cases. Such as when $T\rightarrow\infty$, in this scenario $\rho$ is close to 1, the bound therefore is equivalent to $\mathcal{O}(\frac{2\eta G}{1-q}-\frac{2\eta G}{1-q^{2}}+\frac{2\eta G}{1-q^{B}})$. Fortunately, SGD typically does not need to find a local minima in infinite time, hence $\rho$ is always less than 1.
Appendix D Supplementary Experiments
D.1 Hardware, Software and Source Code
@@
The experiments were conducted on the NVIDIA GeForce RTX 4090. The code were implemented in PyTorch 2.0.0 and leverage the CUDA Toolkit version 11.8. Our comprehensive tests were conducted on AMD EPYC 7763 CPU @1.50GHz with 64 cores under Ubuntu20.04.6 LTS.
[…]
D.2 Additional Verification Experiments
@@
The verification experiments in this section consist of two parts: (1) We first investigate the reasons behind the randomness of previous works under the non-convex setting. We select forgetting samples under different random seeds, and show that previous works exhibit dependence for forgetting samples. This implies that the forgetting effect in prior studies benefits certain samples while negatively impacting others. (2) Second, we perform experiments on a larger-scale dataset and model with more metrics. Prior studies were unable to achieve this due to the costly Hessian computation.
@@
Sensitivity to the forgetting samples. We conduct experiments on a simple CNN and the hyperparameter setup is consistent with the verification experiments I in the main paper. As shown in Figure 4, we provide intuitive insights into the reasons for the decline of prior works in non-convex settings. Specifically, we investigated the $L^{2}$ norm of the approximate parameter change $\|\mathbf{a}^{-U}\|$ and the norm of the exact parameter change across different selections of the forgetting sample. We observed that the performance of NS and IJ depends on the selection of the forgetting data points, i.e., NS and IJ can approximate actual changes well for some data while generating large approximation errors for others under the non-convex setting, as illustrated in Figure 4. The dependency of forgetting data points results in the decrease in distance and correlation coefficient metrics observed in Figure 1 (b). The observed results for NS and IJ are consistent with previous studies on influence functions (Basu et al. (2021; 2020)), where these methods display random behavior under the non-convex setting, resulting in a decrease in correlation with retraining. In contrast, our approach can effectively approximate the actual norm of parameter changes, regardless of the selection of the forgetting sample.
@@
Figure 4: Verification experiments II. Evaluation shows a comparison between the norm of approximate parameter change $\|\mathbf{a}^{-U}\|$ and norm of exact parameter change $\|\mathbf{w}_{E,B}^{-U}-\mathbf{w}_{E,B}\|$ across different random seeds. Intuitively, the NS and IJ methods are contingent on the selection of forgetting data. In contrast, our approach consistently approximates the actual values effectively.
@@
Additional experiments. We further evaluate using larger-scale model ResNet-18 (He et al. (2016)) which features 11M parameters with three datasets: CIFAR-10 (Alex (2009)) for image classification, CelebA (Liu et al. (2015)) for gender prediction, and LFWPeople (Huang et al. (2007)) for face recognition across 29 different individuals. Since previous second-order works (NS Sekhari et al. (2021)) and (IJ Suriyakumar et al. (2022)) are difficult to compute (Hessian) in this scenario, we instead use the following fast baselines, as described and similarly set up in Tarun et al. (2023):
@@
  • FineTune: In the case of finetuning, the learned model is finetuned on the remaining dataset.
@@
  • NegGrad: In the case of gradient ascent, the learned model is finetuned using the negative of update direction on the forgetting dataset.
@@
Configurations: For FineTune and NegGrad, we adjust the hyperparameters and execute the algorithm until the accuracy on the forgetting dataset approaches that of the retrained model. Here are our detailed hyperparameter settings: (1) We conducted evaluation on ResNet-18 trained on CIFAR-10 with 50,000 samples. The learning stage is conducted for 40 epochs with step size of 0.001 and batch size of 256. For FineTune, 10 epochs of training are performed with step size of 0.001 and batch size of 256. For NegGrad, we run gradient ascent for 1 epoch with step size of $5\times 10^{-5}$ and batch size of 2. (2) We conducted evaluation on ResNet-18 trained on LFW with 984 samples, for the classification of 29 facial categories. The learning stage is conducted for 50 epochs with step size of 0.004 and batch size of 40. For FineTune, 10 epochs of training are performed with step size of 0.004 and batch size of 40. For NegGrad, we run gradient ascent for 1 epoch with step size of $4e^{-6}$ and batch size of 1. (3) We conducted evaluation on ResNet-18 trained on CelebA with 10,000 samples. The learning stage is conducted for 5 epochs with step size of 0.01 and batch size of 64. For FineTune, 2 epochs of training are performed with step size of 0.01 and batch size of 64. For NegGrad, we run gradient ascent for 1 epoch with a step size of 0.0005 and batch size of 1. We threshold the gradients within 10 for the above experiments. Finally, we uniformly unlearned 50 data points for each dataset. We denote the test data as $D_{t}$, the remaining data as $D_{r}$, and the forgetting data as $D_{f}$.
@@
The following Table 3-5 are our evaluation results: we employ five widely-used metrics to evaluate the performance of unlearning methods: Acc_$D_{t}$ (accuracy on $D_{t}$), error _$D_{t}$ (error on $D_{t}$), error_$D_{r}$ (error on $D_{r}$), error_$D_{f}$ (error on $D_{f}$), Distance ($L^{2}$ norm between unlearned and retrained model).
@@
Performances of our scheme: We still maintain model indistinguishability from the retrained model on more complex tasks and larger models, where the distance is only 0.06 on CIFAR10 and 0.29 on CelebA. The results also show that the unlearned model of proposed method performs similarly to the retrained model in terms of the accuracy metric. Moreover, compared to these baselines, our method achieves forgetting at only the millisecond level, demonstrating the potential of our method on larger-scale models. Additionally, we noticed that when training is unstable, our method would lead to larger performance degradation, such as on the LFW dataset. Although our method has similar accuracy on the forgetting dataset, it suffers performance loss on the remaining dataset. This implies that our method is sensitive to the training process and hyperparameters in non-convex scenarios. We summarize the limitations in Subsection B.1, and provide possible schemes to address these problems.